OpenStudy (anonymous):

Help for a medal! The sum of two numbers is 0. Twice the smaller number subtracted from 3 times the larger number is 10. Let x represent the larger number and y represent the smaller number. Which equations represent this situation?

OpenStudy (anonymous):

x + y = 0(-2) 2x -3y = 10 ----------- -2x -2y = 0 2x -3y = 10 ----------- -5y = 10 y = -2 if y = -2 x + y = 0 x + (-2) = 0 x = 2

OpenStudy (anonymous):

y=-x y=3/2 is this right?

OpenStudy (anonymous):

y=3/2x-5 i messed up the bottom?

OpenStudy (anonymous):

I'll explain: D

OpenStudy (anonymous):

"The sum of two numbers is 0" as we do not know what these numbers. Let that be super x and y! writing what was said: x + y = 0 (There are two numbers that do not know) "Twice the smallest number subtracted from the most 3 times the apparent 10". Now let's write this second part 2x - 3y = 10 (the exercises said if I multiply by 2 smaller decrease after the second number by multiplying by 3 the result will be 10) then the two equations are thus: x + y = 0 2x - 3y = 10 then just solve the system! : D gave to help? \ o